a bicycle wheel of mass 1.2 kg has a radius of 0.5m . Calculate the moment of inertia of the wheel.
1
Expert's answer
2020-10-05T10:53:36-0400

By defenition, "I=mr^2,"

"I=1.2\space kg \cdot(0.5 m)^2=0.3\space kg\cdot m^2."